# The general Brannan coefficient conjecture and Watson’s lemma

### v1

@@ 2 Integral lower bounds
(9) $A_{n}(\alpha,\beta,\omega)=(-1)^{n+1}\frac{\Gamma(1+\alpha)\Gamma(n-\alpha)\sin(\pi\alpha)}{\pi n!}\\ \times\omega^{n}\ {}_{2}F_{1}\left(\beta,-n;\alpha-n+1;-1/\omega\right).$
@@
for $\alpha,\beta\in(0,1]$, $\phi\in[0,\pi]$ and $n=3,5,7,\ldots$, where
(12) $w_{n}(\alpha,\beta,x)=\sin(\pi\alpha)\,{}_{2}F_{1}\!\left(\beta,-n;\alpha-n+1;x\right).$
Note from Eqs. 1, 4, 9, and 12 that for $n=1,2,3,\ldots$
(13) $\lim_{\alpha\to 0}w_{n}(\alpha,\beta,x)=-\frac{\pi\Gamma(\beta+n)}{\Gamma(\beta)(n-1)!}x^{n},$
@@
(14) $\lim_{\alpha\to 1}w_{n}(\alpha,\beta,x)=\frac{\pi\,\Gamma(\beta+n-1)}{\Gamma(\beta)(n-2)!}\left\{n-(\beta+n-1)x\right\}\,x^{n-1}.$
@@
(15) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}w_{n}(\alpha,\beta,x)\\ =(-1)^{n+1}\sin(\pi\beta)(-x)^{\,n-\alpha}\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,(e^{s}-x)^{\alpha}e^{-ns}\,ds\\ +\sin(\pi\alpha)(-x)^{-\beta}\int_{0}^{\infty}(e^{s}-1)^{\alpha}\left(e^{s}-1/x\right)^{-\beta}e^{-ns}ds.$
@@
From [NIST:DLMF, Eqs. 15.1.2 and 15.6.4]
(16) $_{2}F_{1}(a,b;c;z)=\frac{e^{-b\pi i}\Gamma(1-b)\Gamma(c)}{2\pi i\,\Gamma(c-b)}\int_{1}^{(0+)}\frac{t^{\,b-1}(1-t)^{c-b-1}}{(1-zt)^{a}}\,dt,$
with $b\neq 1,2,3,\dots$, $c\neq 0,-1,-2,\dots$, $\Re(c-b)>0$, and $|\arg(1-z)|<\pi$. Here the contour is the standard loop starting at $t=1$ winding once counterclockwise around $t=0$ and returning to $t=1$.
So with $a=\beta$, b=nb=-n, $c=1+\alpha-n$ and $x=1/\xi$ we have from Eqs. 8 and 12 the integral representation
(17) $w_{n}(\alpha,\beta,1/\xi)=\frac{i\,n!}{2\Gamma(n-\alpha)\Gamma(1+\alpha)}(-\xi)^{\beta}I(\xi),$
@@
(18) $I(\xi)=\int_{1}^{(0+)}\frac{t^{-n-1}(1-t)^{\alpha}}{(t-\xi)^{\beta}}\,dt.$
We temporarily assume $\xi=e^{i\theta}$ with $\theta\in(0,\pi]$ and $\alpha,\beta\in(0,1)$, and for these values the integral converges, $\xi$ lies off the cut $1,\infty)$, and the branch points $t=1$ and $t=\xi$ are distinct. The contour initially lies inside the unit circle $|t|=1$ and so avoids the branch point $t=\xi$. We take the branch cut for $(1-t)^{\alpha}$ to be $1,\infty)$ (principal branch on $\mathbb{C}\setminus[1,\infty)$), and the branch cut for $(t-\xi)^{-\beta}$ to be the ray extending from $\xi$ to $\infty$ with $\arg(t)=\arg(\xi)=\theta$.
Since $\alpha,\beta\in(0,1)$ the branch points at $t=1$ and $t=\xi$ are locally integrable, and as $t\to\infty$ the integrand is $\mathcal{O}(t^{-n-1+\alpha-\beta})$ so contributions from circular arcs vanish at infinity. Hence the loop in Eq. 18 may be deformed to four rays, two on either side of both of the cuts. As a result we arrive at
@@
where $I_{\xi}^{\pm}(\xi)$ denotes the integral of $t^{-n-1}(1-t)^{\alpha}(t-\xi)^{-\beta}$ from $t=\xi$ to $t=\infty\exp(i\theta)$ with $\arg(\xi-t)=\arg(\xi)\mp\pi$, and $I_{1}^{\pm}(\xi)$ denotes the integral of the same integrand from $t=1$ to $t=\infty$ with $\arg(1-t)=\mp\pi$.
@@
Therefore, since $(1-t)^{\alpha}=(1-\xi v)^{\alpha}$ takes the same value for both integrals for $\theta\in(0,\pi)$, we obtain
@@
Inserting Eqs. 20 and 21 into Eq. 19 yields
@@
Now plug Eq. 22 into Eq. 17 and in each of the resulting integrals set $v=e^{s}$, so that $dv=e^{s}\,ds$ and $v^{-n-1}dv=e^{-ns}\,ds$. Then, on using $(1-\xi e^{s})^{\alpha}=(-\xi)^{\alpha}(e^{s}-\xi^{-1})^{\alpha}$ and replacing $\xi$ by $1/x$, we confirm that Eq. 15 holds for $\alpha,\beta\in(0,1)$ and $\arg(1/x)=\theta\in(0,\pi]$. Finally, the identity extends to $\alpha,\beta\in[0,1]$ by continuity (all boundary singularities are removable, with the case $\beta=1$ interpreted by the limiting value of the first term), and it extends from $x=e^{-i\theta}$, $\theta\in(0,\pi]$, to all $x\in\mathbb{C}\setminus[0,\infty)$ by analytic continuation in $x$.
@@
(26) $I_{2}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{\alpha}\left\{\left(1+e^{s}\right)^{-\beta}-R(s,\phi)^{-\beta}\right\}e^{-ns}ds,$
@@
(28) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}}{n!\Delta^{2}}\,\geq h(\alpha,\beta,\phi;n).$
@@
(29) $\frac{\Gamma(n-\alpha)\,\Gamma(1+\alpha)}{n!}\,\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}\\ \geq\sin(\pi\beta)\,\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,\left\{(e^{s}+1)^{\alpha}-\left|e^{s}-e^{i\phi}\right|^{\alpha}\right\}e^{-ns}\,ds\\ \quad+\sin(\pi\alpha)\,\int_{0}^{\infty}(e^{s}-1)^{\alpha}\,\left\{(e^{s}+1)^{-\beta}-\left|e^{s}-e^{-i\phi}\right|^{-\beta}\right\}e^{-ns}\,ds,$
@@ 3 Two-term Watson-type expansions for $I_{1}$ and $I_{2}$
(30) $C_{1}(\alpha,\phi)=\frac{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}{\Delta^{2}},$
(31) $C_{2}(\beta,\phi)=\frac{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}{\Delta^{2}},$
(32) $K_{1}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}-1-\frac{\alpha-\beta}{2}\,s\right],$
@@
(33) $K_{2}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{R(s,\phi)^{-\beta}-(1+e^{s})^{-\beta}}{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}-1-\frac{\alpha-\beta}{2}\,s\right],$
@@
(41) $K_{1}(0,0,\phi;s)=K_{2}(0,0,\phi;s)=\frac{1}{s^{2}}\left[\frac{\ln\left\{\left(1+e^{s}\right)/R(s,\phi)\right\}}{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}-1\right]\quad(0<\phi<\pi).$
@@
The coefficients $C_{1}(\alpha,\phi)$ and $C_{2}(\beta,\phi)$ defined in Eqs. 30 and 31 remain smooth and strictly positive for $\phi\in[\phi_{0},\pi)$ and $\alpha,\beta\in(0,1)$, and have finite nonzero limits as $\phi\to\pi$; in particular
@@
(47) $G_{1}(\alpha,\beta,\phi;s)=\frac{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}}{\Delta^{2}(e^{s}-1)^{\beta}},$
@@
(48) $G_{2}(\alpha,\beta,\phi;s)=\frac{(e^{s}-1)^{\alpha}}{\Delta^{2}}\left[(1+e^{s})^{-\beta}-R(s,\phi)^{-\beta}\right].$
Using a Taylor expansion at $s=0$ and the definition Eqs. 30 and 31 one finds that, as $s\to 0$,
@@
where the remainder kernels $K_{1,2}(\alpha,\beta,\phi;s)$ are given by Eqs. 32 and 33 and satisfy Eq. 38. Substituting Eqs. 49 and 50 into Eqs. 25 and 26, and using
@@
Then, on combining Eqs. 27, 34, 35, 36, 37, 52, and 53, one arrives at:
@@
and in all cases (finite) limits applying when $\alpha\to 0,1$, $\beta\to 0,1$ and/or $\phi\to\pi$. In Eq. 59 $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ are given by Eqs. 24, 32, and $3$.
Note, from Eqs. 52 and 53,
@@
and $L(\alpha,\beta,\pi;s)$ follows from Eqs. 39, 40, 59, 62, and 65.
@@
We also note from Eqs. 23, 24, 32, 33, 59, 61, and 64 that
(69) $L(0,1,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\frac{\csc(\tfrac{1}{2}\phi)-1}{2}\left(1-\frac{s}{2}\right)-\frac{1}{R(s,\phi)}+\frac{1}{1+e^{s}}\right\},$
@@
Before discussing the computations leading to Eqs. 76 and 77, we apply these values to obtain the following desired explicit lower bounds.
@@ 3.1 Numerical computation of $m_{0}$ and $m_{\infty}$
The boundary cases $\alpha=0,1$, $\beta=0,1$, and $\phi=\pi$ are removable singularities; these were evaluated stably using the exact limiting formulas in Eqs. 39, 40, 41, 42, 43, 59, 60, 61, 62, 63, 64, 65, and 73.
@@
$\left(\frac{s}{e^{s}-1}\right)^{\beta}\left\{(1+e^{s})^{\alpha}-R(s,\phi)^{\alpha}\right\}\\ =\left(\frac{s}{1-e^{-s}}\right)^{\beta}e^{(\alpha-\beta)s}\left\{(1+e^{-s})^{\alpha}-\left(1+e^{-2s}-2e^{-s}\cos(\phi)\right)^{1/2}\right\},$
which appears in Eq. 75 via Eqs. 24, 32, and 59. We then minimized the resulting large $s$ stable form of $L_{\infty}(\alpha,\beta,\phi;s)$ over $\alpha,\beta\in(0,1)$, $\phi\in[\phi_{0},\pi)$ and $s\in[30,S]$ using Maple’s global search (Optimization[Minimize]) (using high precision arithmetic), for various values of $S\in[100,10^{5}]$. In every run, the optimizer returned the same minimum at the lower boundary $s=30$, with the value given in Eq. 81. This provides a practical numerical certification that the infimum defining $m_{\infty}$ is achieved within the finite window $1\leq s\leq 30$, bearing in mind, from Eqs. 71 and 75, that $L_{\infty}(\alpha,\beta,\phi;s)=\mathcal{O}(s^{-1})$ as $s\to\infty$. This is illustrated by our optimization computation for $s=10^{5}$ giving the value
@@
Finally, we also used Optimization[Minimize] to independently verify the grid results that yielded Eqs. 76 and $7$ for $s\in[0,1]$ and $s\in[1,30]$, respectively. The Maple source code for the evaluation of $m_{0}$ and $m_{\infty}$ is provided in [Dunster:2026:BWM].

### v2

@@ 2 Integral lower bounds
(9) $A_{n}(\alpha,\beta,\omega)=\frac{(-1)^{n+1}}{\pi}\omega^{n}w_{n}(\alpha,\beta,-1/\omega),$
where
(10) $w_{n}(\alpha,\beta,x)=\frac{1}{n!}\Gamma(n-\alpha)\Gamma(1+\alpha)\sin(\pi\alpha)\,{}_{2}F_{1}\!\left(\beta,-n;\alpha-n+1;x\right).$
@@
for $\alpha,\beta\in(0,1]$, $\phi\in[0,\pi]$ and $n=3,5,7,\ldots$.
Note from Eqs. 1, 4, 9, and 10 that for $n=1,2,3,\ldots$
(13) $\lim_{\alpha\to 0}w_{n}(\alpha,\beta,x)=-\frac{\pi\,\Gamma(\beta+n)}{\Gamma(\beta)\,n!}\,x^{n}.$
@@
(14) $\lim_{\alpha\to 1}w_{n}(\alpha,\beta,x)=\frac{\pi\,\Gamma(\beta+n-1)}{\Gamma(\beta)\,n!}\left\{n-(\beta+n-1)x\right\}\,x^{n-1}.$
@@
(15) $w_{n}(\alpha,\beta,x)=(-1)^{n+1}\sin(\pi\beta)(-x)^{\,n-\alpha}\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,(e^{s}-x)^{\alpha}e^{-ns}\,ds\\ +\sin(\pi\alpha)(-x)^{-\beta}\int_{0}^{\infty}(e^{s}-1)^{\alpha}\left(e^{s}-1/x\right)^{-\beta}e^{-ns}ds.$
@@
From Eq. 1 and Cauchy’s integral formula, for any $c\in(0,1)$ we have for $|\omega|<1$
(16) $A_{n}(\alpha,\beta,\omega)=\frac{1}{2\pi i}\oint_{|z|=c}\frac{(1+\omega z)^{\alpha}}{(1-z)^{\beta}}\,\frac{dz}{z^{n+1}},$
with the contour positively orientated. Setting $\omega=-\xi$ and using Eq. 9 (with $x=1/\xi=-1/\omega$) yields
(17) $w_{n}(\alpha,\beta,1/\xi)=-\frac{1}{2i\xi^{n}}\oint_{|z|=c}\frac{(1-\xi z)^{\alpha}}{(1-z)^{\beta}}\,\frac{dz}{z^{n+1}}.$
Now we temporarily assume $\xi=e^{i\varphi}$ with $\varphi\in(0,\pi]$ and $\alpha,\beta\in(0,1)$, and make the change of variable $z=t/\xi$. As a result we obtain
(18) $w_{n}(\alpha,\beta,1/\xi)=\tfrac{1}{2}i(-\xi)^{\beta}I(\xi),$
@@
(19) $I(\xi)=\oint_{|t|=c}\frac{t^{-n-1}(1-t)^{\alpha}}{(t-\xi)^{\beta}}\,dt,$
with the contour again positively orientated and $c\in(0,1)$.
Next, $\xi=e^{i\varphi}$ lies off the cut $1,\infty)$, and the branch points $t=1$ and $t=\xi$ are distinct. The contour initially lies inside the unit circle $|t|=1$ and so avoids the branch point $t=\xi$. We take the branch cut for $(1-t)^{\alpha}$ to be $1,\infty)$ (principal branch on $\mathbb{C}\setminus[1,\infty)$), and the branch cut for $(t-\xi)^{-\beta}$ to be the ray extending from $\xi$ to $\infty$ with $\arg(t)=\arg(\xi)=\varphi$.
Since we assume that $\alpha,\beta\in(0,1)$, the branch points at $t=1$ and $t=\xi$ are locally integrable, and as $t\to\infty$ the integrand is $\mathcal{O}(t^{-n-1+\alpha-\beta})$, so contributions from circular arcs vanish at infinity. Hence, the loop in Eq. 19 may be deformed to four rays, two on either side of both cuts. Consequently, we arrive at
@@
where $I_{\xi}^{\pm}(\xi)$ denotes the integral of $t^{-n-1}(1-t)^{\alpha}(t-\xi)^{-\beta}$ from $t=\xi$ to $t=\infty\exp(i\varphi)$ with $\arg(\xi-t)=\arg(\xi)\mp\pi$, and $I_{1}^{\pm}(\xi)$ denotes the integral of the same integrand from $t=1$ to $t=\infty$ with $\arg(1-t)=\mp\pi$.
@@
Therefore, since $(1-t)^{\alpha}=(1-\xi v)^{\alpha}$ takes the same value for both integrals for $\varphi\in(0,\pi)$, we obtain
@@
Inserting Eqs. 21 and 22 into Eq. 20 yields
@@
Now plug Eq. 23 into Eq. 18 and in each of the resulting integrals set $v=e^{s}$, so that $dv=e^{s}\,ds$ and $v^{-n-1}dv=e^{-ns}\,ds$. Then, on using $(1-\xi e^{s})^{\alpha}=(-\xi)^{\alpha}(e^{s}-\xi^{-1})^{\alpha}$ and replacing $\xi$ by $1/x$, we confirm that Eq. 15 holds for $\alpha,\beta\in(0,1)$ and $\arg(1/x)=\varphi\in(0,\pi]$. Finally, the identity extends to $\alpha,\beta\in[0,1]$ by continuity (all boundary singularities are removable, with the case $\beta=1$ interpreted by the limiting value of the first term), and it extends from $x=e^{-i\varphi}$, $\varphi\in(0,\pi]$, to all $x\in\mathbb{C}\setminus[0,\infty)$ by analytic continuation in $x$.
@@
(27) $I_{2}(\alpha,\beta,\phi;n)=\frac{1}{\Delta^{2}}\int_{0}^{\infty}\left(e^{s}-1\right)^{\alpha}\left\{\left(e^{s}+1\right)^{-\beta}-R(s,\phi)^{-\beta}\right\}e^{-ns}ds,$
@@
(29) $\Delta^{-2}\left\{w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\right\}\geq h(\alpha,\beta,\phi;n).$
@@
(30) $w_{n}(\alpha,\beta,-1)-\left|w_{n}(\alpha,\beta,e^{i\phi})\right|\\ \geq\sin(\pi\beta)\,\int_{0}^{\infty}(e^{s}-1)^{-\beta}\,\left\{(e^{s}+1)^{\alpha}-\left|e^{s}-e^{i\phi}\right|^{\alpha}\right\}e^{-ns}\,ds\\ \quad+\sin(\pi\alpha)\,\int_{0}^{\infty}(e^{s}-1)^{\alpha}\,\left\{(e^{s}+1)^{-\beta}-\left|e^{s}-e^{-i\phi}\right|^{-\beta}\right\}e^{-ns}\,ds,$
@@ 3 Two-term Watson-type expansions for $I_{1}$ and $I_{2}$
(31) $C_{1}(\alpha,\phi)=\Delta^{-2}\left[2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}\right],$
(32) $C_{2}(\beta,\phi)=\Delta^{-2}\left[\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}\right],$
(33) $K_{1}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{s}{e^{s}-1}\right)^{\beta}\frac{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}}{2^{\alpha}-\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{\alpha}}-1-\frac{\alpha-\beta}{2}\,s\right],$
@@
(34) $K_{2}(\alpha,\beta,\phi;s)=\frac{1}{s^{2}}\left[\left(\frac{e^{s}-1}{s}\right)^{\alpha}\frac{R(s,\phi)^{-\beta}-(e^{s}+1)^{-\beta}}{\left\{2\sin(\tfrac{1}{2}\phi)\right\}^{-\beta}-2^{-\beta}}-1-\frac{\alpha-\beta}{2}\,s\right],$
@@
(42) $K_{1}(0,0,\phi;s)=K_{2}(0,0,\phi;s)=\frac{1}{s^{2}}\left[\frac{\ln\left\{\left(e^{s}+1\right)/R(s,\phi)\right\}}{\ln\left\{\csc(\tfrac{1}{2}\phi)\right\}}-1\right]\quad(0<\phi<\pi).$
@@
The coefficients $C_{1}(\alpha,\phi)$ and $C_{2}(\beta,\phi)$ defined in Eqs. 31 and 32 remain smooth and strictly positive for $\phi\in[\phi_{0},\pi)$ and $\alpha,\beta\in(0,1)$, and have finite nonzero limits as $\phi\to\pi$; in particular
@@
(48) $G_{1}(\alpha,\beta,\phi;s)=\frac{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}}{\Delta^{2}(e^{s}-1)^{\beta}},$
@@
(49) $G_{2}(\alpha,\beta,\phi;s)=\frac{(e^{s}-1)^{\alpha}}{\Delta^{2}}\left[(e^{s}+1)^{-\beta}-R(s,\phi)^{-\beta}\right].$
Using a Taylor expansion at $s=0$ and the definition Eqs. 31 and 32 one finds that, as $s\to 0$,
@@
where the remainder kernels $K_{1,2}(\alpha,\beta,\phi;s)$ are given by Eqs. 33 and 34 and satisfy Eq. 39. Substituting Eqs. 50 and 51 into Eqs. 26 and 27, and using
@@
Then, on combining Eqs. 28, 35, 36, 37, 38, 53, and 54, one arrives at:
@@
and in all cases (finite) limits applying when $\alpha\to 0,1$, $\beta\to 0,1$ and/or $\phi\to\pi$. In Eq. 60 $K_{1}(\alpha,\beta,\phi;s)$ and $K_{2}(\alpha,\beta,\phi;s)$ are given by Eqs. 25, 33, and 34.
Note, from Eqs. 53 and 54,
@@
and $L(\alpha,\beta,\pi;s)$ follows from Eqs. 40, 41, 60, 63, and $6$.
@@
We also note from Eqs. 24, 25, 33, 34, 60, 62, and 65 that
(70) $L(0,1,\phi;s)=\frac{1}{\pi(\pi-\phi)^{2}}\left\{\frac{\csc(\tfrac{1}{2}\phi)-1}{2}\left(1-\frac{s}{2}\right)-\frac{1}{R(s,\phi)}+\frac{1}{e^{s}+1}\right\},$
@@
Before discussing the computations leading to Eqs. 77 and 78, we apply these values to obtain the following desired explicit lower bounds.
@@ 3.1 Numerical computation of $m_{0}$ and $m_{\infty}$
The boundary cases $\alpha=0,1$, $\beta=0,1$, and $\phi=\pi$ are removable singularities; these were evaluated stably using the exact limiting formulas in Eqs. 40, 41, 42, 43, 44, 60, 61, 62, 63, 64, 65, 66, and 74.
@@
$\left(\frac{s}{e^{s}-1}\right)^{\beta}\left\{(e^{s}+1)^{\alpha}-R(s,\phi)^{\alpha}\right\}\\ =\left(\frac{s}{1-e^{-s}}\right)^{\beta}e^{(\alpha-\beta)s}\left\{(1+e^{-s})^{\alpha}-\left(1+e^{-2s}-2e^{-s}\cos(\phi)\right)^{1/2}\right\},$
which appears in Eq. 76 via Eqs. 25, 33, and 60. We then minimized the resulting large $s$ stable form of $L_{\infty}(\alpha,\beta,\phi;s)$ over $\alpha,\beta\in(0,1)$, $\phi\in[\phi_{0},\pi)$ and $s\in[30,S]$ using Maple’s global search (Optimization[Minimize]) (using high precision arithmetic), for various values of $S\in[100,10^{5}]$. In every run, the optimizer returned the same minimum at the lower boundary $s=30$, with the value given in Eq. 82. This provides a practical numerical certification that the infimum defining $m_{\infty}$ is achieved within the finite window $1\leq s\leq 30$, bearing in mind, from Eqs. 72 and 76, that $L_{\infty}(\alpha,\beta,\phi;s)=\mathcal{O}(s^{-1})$ as $s\to\infty$. This is illustrated by our optimization computation for $s=10^{5}$ giving the value
@@
Finally, we also used Optimization[Minimize] to independently verify the grid results that yielded Eqs. 77 and 78 for $s\in[0,1]$ and $s\in[1,30]$, respectively. The Maple source code for the evaluation of $m_{0}$ and $m_{\infty}$ is provided in [Dunster:2026:BWM].